OpenStudy (anonymous):

What is the quotient: (6x4 - 15x3 + 10x2 - 10x + 4) ÷ (3x2 + 2) ? Choose one answer. a. 2x2 - 5x + 2 b. 2x2 + 5x - 2 c. 2x2 - 5x - 2 d. 2x2 + 5x + 2

OpenStudy (anonymous):

Solve for x over the real numbers: 2 x^2-5 x+2 = 0 Factor the left hand side. The left hand side factors into a product with two terms: (x-2) (2 x-1) = 0 Solve each term in the product separately. Split into two equations: x-2 = 0 or 2 x-1 = 0 Look at the first equation: Solve for x. Add 2 to both sides: x = 2 or 2 x-1 = 0 Look at the second equation: Isolate terms with x to the left hand side. Add 1 to both sides: x = 2 or 2 x = 1 Solve for x. Divide both sides by 2: Answer: | | x = 2 or x = 1/2